# A Clinically-Grounded Two-Stage Framework for Renal CT Report Generation

### v1

@@ 6 Limitations
Our dataset is relatively small and exhibits substantial class imbalance, particularly for rare feature labels such as endophytic and tumor. For example, the seemingly perfect performance reported for the Exophytic category is due to a single minority instance appearing in both the training and validation sets, resulting in inflated metrics. Expanding the dataset and improving class balance will be essential for obtaining more reliable performance estimates [9, 10].
@@
Limitations of Text-Based Evaluation
The reference sentences used for evaluating report generation were extracted from clinician-written reports that vary in style, wording, and completeness. Traditional NLG metrics such as BLEU and ROUGE are limited in their ability to assess clinical accuracy.
@@ 7 Conclusion
This study presents a clinically grounded framework for generating renal radiology reports from 2D CT slices, combining structured abnormality detection with multimodal report generation. Our approach mirrors key steps in the radiology workflow and demonstrates the feasibility of integrating feature-aware modeling with large vision-language models in a modular and interpretable manner.
While the current system is limited to 2D slices, it lays the groundwork for more comprehensive and clinically realistic solutions. In future work, we plan to extend the framework to full 3D CT volumes to capture richer anatomical context and enable more robust report generation. Such advancements may further support clinical decision-making in renal cancer diagnosis and management.

### v2

@@ 5.3 Limitations
Our dataset is relatively small and exhibits substantial class imbalance, particularly for rare feature labels such as endophytic and tumor. For example, the seemingly perfect performance on the Exophytic category is due to a single minority instance appearing in both training and validation sets, resulting in inflated metrics. In addition, the rarity of the endophytic label may reflect reporting conventions rather than true clinical frequency, as radiologists at our institution may not consistently use this descriptor. Future work should incorporate multi-center data annotated by diverse radiologists to better capture reporting heterogeneity and improve generalizability. Expanding the dataset and improving class balance will be essential for obtaining more reliable performance estimates [13, 14].
@@
Non-uniform Slice Thickness
In this work, we did not standardize the slice thickness of CT scans in physical space. This variability may lead to inconsistent spatial coverage during data augmentation, which can in turn affect lesion size estimation and the model’s spatial reasoning. Future work should incorporate slice spacing metadata from DICOM headers or resample all scans to a uniform voxel resolution to ensure consistent 3D context.
@@
Loss of Clinical Feature Fidelity in Generated Reports
While the generated reports often appear fluent and well-structured, they may fail to accurately represent the intended clinical features, particularly in fine-tuned models. This discrepancy becomes evident during automatic clinical feature extraction, where key findings are occasionally omitted or hallucinated (e.g., incorrect lesion size, false tumor attribution). Such deviations reduce the clinical reliability of the generated reports and underscore a gap between surface-level textual quality and underlying clinical fidelity.
@@ 6 Conclusion
This study presents a clinically informed, two-stage framework for renal radiology report generation from 2D CT slices. The first stage detects structured clinical features, and the second stage generates corresponding free-text reports conditioned on both the CT image and the detected features. This design enables the model to simultaneously produce structured and unstructured outputs, closely aligning with the dual needs of radiology workflows—supporting both quantitative feature analysis and narrative documentation.
While the current system is limited to 2D slices, it lays the groundwork for more comprehensive and clinically realistic solutions. In future work, we plan to extend the framework to full 3D CT volumes to capture richer anatomical context and enable more robust report generation. Additional efforts may focus on improving the clinical fidelity of generated reports through fact-checking modules, constrained decoding, or feature-aware training objectives. Expanding the dataset and improving label completeness—potentially through multi-source annotation or expert refinement—will also be critical for addressing data sparsity and class imbalance.
Funding
Not available.
Author contributions
RL: Conceptualization, Data curation, Methodology, Writing – original draft. ZF: Conceptualization, Data curation. JP: Methodology, Validation. CS: Methodology, Validation. RT: Formal analysis, Supervision, Writing – review & editing. BS: Data curation. JX: Conceptualization, Supervision, Writing – review & editing.
Competing interests
The authors declare that they have no competing interests.
Data availability
Clinical CT images and radiology reports were obtained from the UF Health Integrated Data Repository (IRB202400720). Due to HIPAA regulations and institutional data-use agreements, the raw clinical data cannot be shared publicly. To support reproducibility, we provide the full preprocessing and training code at https://github.com/renjie-liang/two-stage-renal-ct-report-generation, along with a small synthetic dataset that follows the same structure and labeling schema as the real data.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this manuscript, the authors used ChatGPT to improve the readability and language. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article. No patient data or protected health information (PHI) were provided to the tool.